please solve the initial value problem
3-04 +2=– 1 - 0 3, 39 = 0, v9-5 y – 3y + 2y = 58(t – 1) - U2 (t)e-2, y(0) = 0, y (0) = 5
Answer #1

t-2 Given that the initial value problem yl-34 +2y = 58(E-) - Ug (t) er .. ① y (0)20, y (0-5 Here &(t-D is a dirac delta fun:. 12 A1 (5-284) + B1 (8 35+2) + G (S-2). = (A1+B1) S + (-2A1-3 Bj+G) St A1 + 2B, 29. in AltB120 B, 2-AI B1D-11. ge) - Ilford